### v1

# Bruhat operads

Abstract
We describe some planar operads build from the higher Bruhat orders.
M. Jourdan: Par ma foi, il y a plus de quarante ans
que je dis de la prose sans que j’en susse rien
Bourgeois gentilhomme
1 Introduction
1.1
The aim of the present paper is to introduce some planar operads build from higher Bruhat orders, cf. [2].
Recall that a planar111sometimes called nonsymmetric operad of sets $\mathcal{O}$ consists of a family of sets $\{\mathcal{O}(n)\}_{n\in\mathbb{Z},n\geq 1}$, together with composition maps
$\gamma:\ \mathcal{O}(n)\times\mathcal{O}(n_{1})\times\ldots\times\mathcal{O}(n_{m})\to\mathcal{O}(\sum_{i}n_{i})$
and a unit $1\in\mathcal{O}(1)$ satisfying the associativity and the unit axiom as for a usual operad, except that we do not require an action of symmetric groups $S(n)$ on $\mathcal{O}(n)$, cf. [5], Def. 3.12 and Def. 2.1 below.
1.2
Let $d,n$ be integers, $1\leq d\leq n$. In [2] (cf. also [3]) certain posets $B(n,d)$ called higher Bruhat orders have been introduced, their definition is recalled below in 5.1.
One can describe them roughly as follows. $B(n,1)=S(n)$ with the classical weak Bruhat order. The poset $S(n)$ admits unique maximal and minimal elements, namely, the minimal element is the unit permutation $e$, whereas the maximal one is the permutation $\sigma_{max}$ of maximal length222where length means the usual length $\ell$ in a Coxeter group, so $\ell(\sigma_{max})=n/(n-1)/2$.. By definition $B(n,1)$ is certain quotient of the set of maximal (i.e. noncompressible) chains in $S(n)$ connecting $e$ with $\sigma_{max}$.
Afterwards one proceeds by induction on $d$. Each $B(n,d)$ is a poset admitting unique minimal and maximal elements, say $b_{min}^{d}$ and $b_{max}^{d}$, and $B(n,d+1)$ is certain quotient of the set of maximal chains in $B(n,d)$ connecting $b_{min}^{d}$ with $b_{max}^{d}$.
A remarkable paper by Ziegler [6] contains a number of deep constructions and results on higher Bruhat orders. They are essential in our study.
1.3
Fix a positive integer $d$. Denote $\mathcal{O}(n):=B(nd,d)$. The first main result of the present paper, see Thm 6.1 below, says that the collection $\mathcal{O}=\{\mathcal{O}(n)\}$ admits a structure of a planar operad, to be called the small Bruhat operad.
Our definition of multiplications $\gamma$ is based on the following main construction.
Given two elements $b\in B(n,d)$ and $b^{\prime}\in B(m,d)$ and an integer $j$, $0\leq j\leq n+m-d$, we define a new element $b\circ_{j}b^{\prime}\in B(n+m-d)$, see 5.6.
The definition of the operations $\circ_{j}$ is nontrivial and uses an important theorem by Ziegler, see Thm. 5.1 below.
1.4
We may include into the game elements of $B(m,d)$ with $d$ not dividing $m$ as well. Namely, we define a planar operad $\mathcal{O}^{\prime}=\{\mathcal{O}^{\prime}(n)\}$ which contains $\mathcal{O}$ as a suboperad and called the big Bruhat operad. The sets $\mathcal{O}^{\prime}(n)$ are certain infinite unions of sets $B(m,d)$ with fixed $d$ and arbitrary $m\geq d$.
This is the second main result of the present paper, see 6.2.
1.5
The contents of the paper is as follows. In Section 2 we recall the definition of planar operads. In Section 3 we discuss certain ”Master planar operads” which serve as a combinatorial base for constructing the big Bruhat operad. In Section 4 we discuss some operads related to symmetric groups which correspond to the case $d=1$ of our considerations. Section 5 contains the main constructions.
1.6 Notation
For $n\in\mathbb{Z},n\geq 1$, $[n]:=\{1,\ldots,n\}$.
For $a\in\mathbb{R},a+[n]:=\{a+1,\ldots a+n\}$.
By (${[n]}\choose d$ we denote the set of $d$ element subsets of $[n]$.
2 Planar operads
2.1
Let $K$ be a monoidal category. Below we will be interested mostly in the category of sets with the monoidal structure given by the cartesian product.
Definition 2.1
A planar operad in $K$ consists of objects $\mathcal{P}(n)$ of $K$ indexed by the natural numbers $n\in{\mathbb{Z}},n\geq 1$, with operations of composition
$\gamma_{k;n_{1},\ldots,n_{k}}\,\colon\,\mathcal{P}(k)\otimes\mathcal{P}(n_{1})\otimes\cdots\otimes\mathcal{P}(n_{k})\to\mathcal{P}(n_{1}+\cdots+n_{k})$ (2.1)
The composition should be associative which means that
$\gamma_{k;n_{1},\ldots,n_{k}}\prod_{1}^{k}\gamma_{n_{t};l_{m_{1}}^{t},\ldots,l_{m_{n_{t}}}^{t}}=\gamma_{k;\sum_{j=1}^{n_{1}}l^{1}_{m_{j}},\ldots,\sum_{j=1}^{n_{k}}l^{k}_{m_{j}}}$ (2.2)
and should have a unit element $I\in P(1)$ such that
$\gamma_{k,I,\ldots,I}=\text{Id}_{\mathcal{P}(k)}.$
2.2 Partial compositions, or insertions.
To define compositions (2.1) it is sufficient to define their particular cases, ”insertions”
$\circ_{i}:\mathcal{P}(m)\otimes\mathcal{P}(n)\to\mathcal{P}(m+n-1)$ (2.3)
where
$a\circ_{i}b=\gamma_{m;1,\ldots,1,n,1,\ldots,1}(a;I,\ldots,I,b,I,\ldots,I),$
$1\leq i\leq m$, satisfying two identities:
for all $a\in\mathcal{P}(l),b\in\mathcal{P}(m),c\in\mathcal{P}(n)$ we should have
$(a\circ_{i}b)\circ_{i-1+j}c=a\circ_{i}(b\circ_{j}c),$
$1\leq i\leq l,1\leq j\leq m$ (associativity), and
$(a\circ_{i}b)\circ_{k-1+m}c=(a\circ_{k}b)\circ_{i}c,$
$1\leq i<k\leq l$ (commutativity), see [1], 5.9.4.
This criterion will be crucial for us.
3 Master planar operads
3.1 Operad $M$.
Consider the following collection of sets $M(n)$, $n\geq 1$.
The set $M(n)$ is consists of collections of integer numbers $\mathbf{k}:=(k_{0},k_{1},\ldots,k_{n})$, $k_{i}\geq 0$.
We can imagine elements of $M(n)$ as collections of black and white different points on the real line such that first (from left to right) we see $k_{0}$ white points, then a black point, then $k_{1}$ white points, then the second black point, then $k_{2}$ white points, and so on until the $n$-th black point and $k_{n}$ white points.
Let us define a family of maps
$\gamma_{n;l_{1},\cdots,l_{n}}:\ M(n)\times M(l_{1})\times\cdots\times M(l_{n})\to M(\sum_{j=1}^{n}l_{j})$ (3.1)
by the following rule.
For $\mathbf{k}:=(k_{0},k_{1},\ldots,k_{n})\in M(n)$, $\mathbf{m}_{1}:=(m_{0}^{1},m_{1}^{1},\ldots,m_{l_{1}}^{1})\in M(l_{1}),\cdots$, $\mathbf{m}_{n}:=(m_{0}^{n},m_{1}^{n},\ldots,m_{l_{n}}^{n})\in M(l_{n})$, we define
$\gamma_{n;l_{1},\cdots,l_{n}}(\mathbf{k},\mathbf{m}_{1},\cdots,\mathbf{m}_{n})=$
$(k_{0}+m_{0}^{1},m_{1}^{1},\cdots,m^{1}_{l_{1}-1},k_{1}+m_{l_{1}}^{1}+m_{0}^{2},m_{1}^{2},\cdots,m_{l_{2}-1}^{2},k_{2}+m_{l_{2}}^{2}+m_{0}^{3},\cdots,$
$k_{n-1}+m_{l_{n-1}}^{n-1}+m_{0}^{n},m_{1}^{n},\cdots,m_{l_{n}-1}^{n},k_{n}+m_{l_{n}}^{n}).$ (3.2)
Geometrically the composition $\gamma$ means that we insert the whole collection $M(1)$ instead of the first black point in $\mathbf{k}$, $M(2)$ instead of the second black point in $\mathbf{k}$, $\ldots$, $M(n)$ instead of the $n$-th black point in $\mathbf{k}$, see the picture below.
The unit element is the tuple $I=(0,0)\in M(1)$.
It is easy to check that the map (3.1) is associative and $I$ is the unit. Thus we have
Lemma 3.1
The collection $M=\{M(n),n\geq 1\}$, with the family of maps (3.1) is a planar operad of sets.
$\square$
Note that the sets $M(n)$ are infinite.
3.2 Operads of molecules $M^{d}$.
For a given positive integer $d$, and $\mathbf{k}=(k_{0},\cdots,k_{n})\in M(n)$, consider the collection
$\mathbf{k}^{d}:=(k_{0},d,k_{1},d,k_{2},d,\cdots,k_{n-1},d,k_{n}).$
The sum of these numbers will be denoted by $|\mathbf{k}^{d}|$; it is equal to $dn+\sum_{j}\ k_{j}$.
A sequence $\mathbf{k^{d}}$ will be called a type.
Let us consider the set $M(\mathbf{k}^{d})=[dn+\sum_{j}k_{j}]$. Its elements will be called particles.
Subsets of $M(\mathbf{k}^{d})$ having the form of intervals
$k_{j}+dj+[d]=\{k_{j}+dj+1,\cdots,k_{j}+d(j+1)\}$
we will call teams, or nuclei, $j=0,\cdots,n-1$. Elements of nuclei will be called protons; so each nucleus contains $d$ protons.
Particles of $M(\mathbf{k}^{d})$ which do not belong to any nucleus will be called idles, or electrons.
The whole set $M(\mathbf{k}^{d})$ equipped with the collection of $n$ nuclei inside it will be called a molecule of type $\mathbf{k}^{d}$. Thus $M(\mathbf{k}^{d})$ contains $|\mathbf{k}^{d}|$ particles, among them $nd$ protons and $\sum_{j}k_{j}$ electrons.
Obviously the type defines the molecule uniquely.
Denote by $M^{d}(n)$ the set of all molecules with $n$ nuclei and arbitrary number of electrons. We can identify $M^{1}(n)$ with $M(n)$, with nuclei (resp. electrons) corresponding to black (resp. white) points.
Define a collection of maps
$\gamma_{n;l_{1},\ldots,l_{n}}:\ M^{d}(n)\times M^{d}(l_{1})\times\cdots\times M^{d}(l_{n})\to M^{d}(\sum_{j=1}^{n}l_{j})$ (3.3)
For molecules $M(\mathbf{k})\in M^{d}(n)$, $\mathbf{k}=(k_{0},d,k_{1},d,\ldots,d,k_{n})$, $M(\mathbf{m}_{1})\in M^{d}(l_{1})$, $\mathbf{m}_{1}=(m_{0}^{1},d,m_{1}^{1},d,\ldots,d,m_{l_{1}}^{1}),\ldots$, $M(\mathbf{m}_{n})\in M^{d}(l_{n})$, $\mathbf{m}_{n}=(m_{0}^{n},d,m_{1}^{n},d,\ldots,d,m_{l_{n}}^{n})$, we set
$\gamma_{n;l_{1},\ldots,l_{n}}(M(\mathbf{k}),M(\mathbf{m}_{1}),\ldots,M(\mathbf{m}_{n}))=$
$M(k_{0}+m_{0}^{1},d,m_{1}^{1},d,\ldots,m^{1}_{l_{1}-1},d,k_{1}+m_{l_{1}}^{1}+m_{0}^{2},d,m_{1}^{2},\ldots,m_{l_{2}-1}^{2},d,k_{2}+m_{l_{2}}^{2}+m_{0}^{3},d,\ldots,$
$d,k_{n-1}+m_{l_{n-1}}^{n-1}+m_{0}^{n},d,m_{1}^{n},\ldots,d,m_{l_{n}-1}^{n},d,k_{n}+m_{l_{n}}^{n})).$ (3.4)
This operation means that instead of the $j$-th nucleus in $M(\mathbf{k})$ we insert the molecule $M(\mathbf{m}_{j})$, $1\leq j\leq n$.
Lemma 3.2
The collection $M^{d}=\{M^{d}(n),n\geq 1\}$, with the family of maps (3.4) is a planar operad isomorphic to $M^{1}$.
Proof Obvious. $\Box$
Note that the sets $M^{d}(n)$ are infinite.
4 Symmetric groups
4.1 A symmetric operad $S$.
It is wellknown that the symmetric groups form an operad.
Namely, let $S(n)$ be the group of permutations of the set $[n]=\{1,\cdots,n\}$.
The family of maps
$\gamma_{k,n_{1},\cdots,n_{k}}:S(k)\times S(n_{1})\times\ldots\times S(n_{k})\to S(n_{1}+\ldots+n_{k})$ (4.1)
defined by sending
$(\tau,\sigma_{1},\ldots\sigma_{k}),\quad\tau\in S(k),\,\sigma_{i}\in S(n_{i}),\,i=1,\cdots,k,$
to the permutation of the set $[n_{1}+\cdots+n_{k}]$ obtained by applying $\sigma_{i}$ to the $i$th interval $n_{1}+\cdots+n_{i-1}+1,\ldots,n_{1}+\cdots+n_{i-1}+n_{i}$, and then permuting these intervals following $\tau$.
The collection $S=\{S(n),n\geq 1\}$, with the above family of maps $\gamma$ is a planar operad, and even more it is a symmetric operad.
4.2 A planar operad $F$.
We introduce another planar operad $F$ on symmetric groups which is related to the master operad $M^{1}$.
We define $F(n)$ as a subset of the set of permutations of $M^{1}(n)$.
Specifically, we consider collections
$\mathbf{k}=(k_{0},1,k_{1},1,k_{2},\cdots,k_{n-1},1,k_{n})\in M^{1}(n)$
with teams of size $1$.
For any set $[n+\sum_{i=0}^{n}k_{i}]$ we consider permutations $\sigma\in S(n+\sum_{i}k_{i})$ which permutes only teams and are identical on the idles.
The set $F(n)$ consists of all such permutations for all $\mathbf{k}\in M^{1}(n)$.
Compositions
$\gamma:F(k)\times F(n_{1})\times\ldots\times F(n_{k})\to F(n_{1}+\ldots+n_{k})$
are defined as follows.
For $\tau\in F(k)$, $\sigma_{i}\in F(n_{i})$, $i=1,\ldots,k$,
$\gamma(\tau;\sigma_{1},\ldots,\sigma_{k})$
permutes
$\gamma(\mathbf{k};\mathbf{m}_{1},\ldots,\mathbf{m}_{k}),\quad\mathbf{k}\in M^{1}(k),\,\mathbf{m}_{i}\in M^{1}(n_{i}),i=1,\ldots,k,$
by permuting the blocks $\mathbf{m}_{i}$ by $\sigma_{i}$ and then permuting these blocks by $\tau$.
We have
Proposition 4.1
The collection of $F(n)$, $n=1,\cdots$ with the above defined family of maps is a planar operad.
Proof follows from Lemma 3.2 for $d=1$. $\Box$
Remark 4.2
Restricting the operad $F(n)$ to the subset of collections without idles gives us the operad $S$ on symmetric groups $S(n)$, $n=1,\ldots$.
4.3 Weak Bruhat order
On a symmetric group $S(n)$, there is a weak Bruhat order $\prec$ which may be defined as follows: $u\prec v$ iff the sets of inversions are nested, $Inv(u)\subset Inv(v)$.
Both operads $S$ and $F$ are compatible with Bruhat orders. Namely
Theorem 4.3
Let $\tau,\tau^{\prime}\in S(k)$ (resp. $\tau,\tau^{\prime}\in F(k)$), $\sigma_{i},\sigma^{\prime}_{i}\in S(n_{i})$ (resp. $\sigma_{i},\sigma^{\prime}_{i}\in F(n_{i})$), $i=1,\ldots k$. If $\tau\prec\tau^{\prime}$ and $\sigma_{i}\prec\sigma^{\prime}_{i}$, $i=1,\ldots,k$, then
$\gamma(\tau;\sigma_{1},\ldots,\sigma_{k})\prec\gamma(\tau^{\prime};\sigma_{1}^{\prime},\ldots,\sigma_{k}^{\prime}).$
We prove this theorem below in Section 5, see 5.8.
5 Higher Bruhat orders and insertions
5.1 Higher Bruhat orders.
Let us recall some definitions from [2].
Let $d\leq n$ be two positive integers. Let $Lin(n,d)$ denote the set of linear orders on the discrete Grassmannian (${[n]}\choose d$.
For a subset $\mathbf{i}=\{i_{1},\ldots,i_{d+1}\}\in{{[n]}\choose d+1}$, $1\leq i_{1}<\cdots<i_{d+1}\leq n$, and $11\leq k\leq d+$ 1 let us denote $\widehat{\mathbf{i}}_{k}$ the sequence $\mathbf{i}$ with $i_{k}$ omitted.
Let us call its packet, and denote by $P(\mathbf{i})$, the sequence
$P(\mathbf{i})=(\widehat{\mathbf{i}}_{d+1},\widehat{\mathbf{i}}_{d},\ldots,\widehat{\mathbf{i}}_{1}).$
Let us call an element $\prec\in L(n,d)$ admissible if its restriction to each packet $P(\mathbf{i})$ is either lexicographical, i.e.
$\widehat{\mathbf{i}}_{d+1}\prec\widehat{\mathbf{i}}_{d}\prec\ldots\prec\widehat{\mathbf{i}}_{1}$
or antilexicographical, i.e.
$\widehat{\mathbf{i}}_{d+1}\succ\widehat{\mathbf{i}}_{d}\succ\ldots\succ\widehat{\mathbf{i}}_{1}.$
Let $A(n,d)\subset L(n,d)$ denote the subset of admissible linear orders.
The set $B(n,d)$ of Bruhat orders for (${[n]}\choose d$ is defined as the quotient $A(n,d)/\sim$ where $\sim$ is an equivalence relation described in [2], no 2.
Each $B(n,d)$ is a poset, to be denoted $\mathcal{HB}^{d}(n)$. One of main results of [2] states that any order from $\mathcal{HB}^{d}(n)$ can be included into a maximal chain joining the minimal order with the maximal one.
5.2 Example
$A(n,1)=B(n,1)=S(n)$ with the classical weak Bruhat order.
5.3 Remark
Alternatively $B(n,d)$ may be defined as the set of minimal partial orders on (${[n]}\choose d$ whose restriction to each packet is either lexicographical or antilexocographical.
5.4 Inversion sets
Let $a\in A(n,d)$. The collection of all $\mathbf{i}\in$ (${[n]}\choose{d+1}$ such that $P(\mathbf{i})$ is ordered anilexicographically forms the set of inversions $Inv(a)\subset$ (${[n]}\choose{d+1}$. If $a\sim a^{\prime}$ then $Inv(a)=Inv(a^{\prime})$, so for $\prec\in B(n,d)$ the set $Inv(\prec)\subset$ (${[n]}\choose{d+1}$ is well defined.
The set of inversions determines the order uniquely.
5.5 Ziegler criterion
Ziegler [6], Th. 4.1(B), characterised the subsets of (${[n]}\choose{d+1}$ which appear as sets of inversions of Bruhat orders.
Theorem 5.1
(Ziegler) A subset $I\subset{{[n]}\choose{d+1}}$ is the set of inversions $Inv(\prec)$ of a Bruhat order $\prec$ for (${[n]}\choose d$ if and only if for any $\mathbf{i}\in{{[n]}\choose{d+2}}$ the intersection of $I$ and the packet $P(\mathbf{i})$ is either a beginning or an ending interval of $P(\mathbf{i})$.
The last condition means that there exists $k$, $20\leq k\leq d+$ 2 such that $P(\mathbf{i})\cap I$ is equal to $(\widehat{\mathbf{i}}_{k},\widehat{\mathbf{i}}_{k-1},\ldots,\widehat{\mathbf{i}}_{1})$ or to $(\widehat{\mathbf{i}}_{d+2},\widehat{\mathbf{i}}_{d+1},\ldots,\widehat{\mathbf{i}}_{k})$. Here $k=0$ (resp. $k=d+2$) means that $P(\mathbf{i})\cap I=\emptyset$ (resp. $P(\mathbf{i})\cap I=P(\mathbf{i})$).
We will say that $\mathbf{i}$ satisfies the Ziegler criterion, or simply that $\mathbf{i}$ satisfies (Z), for $I$.
Let us denote by $Z(n,d+1)$ the set of subsets $I\subset{{[n]}\choose{d+1}}$ such that for any $\mathbf{i}\in{{[n]}\choose{d+2}}$, $\mathbf{i}$ satisfies (Z) for $I$. The above statements mean that $Inv$ establishes a bijection
$Inv:\ B(n,d)\overset{\sim}{\to}Z(n,d+1).$
Note that a maximal chain in the poset $\mathcal{HB}^{d}(n)$ has the following chain of inversion sets. It starts with the empty set, and each consequent one is obtained by adding one more inversion. The length of such a chain is ($[n]\choose d+1$, the cardinality of the inversion set for the maximal element of $\mathcal{HB}^{d}(n)$.
5.6 Main construction: insertions of higher Bruhat orders
To define an operad structure on higher Bruhat orders we use the following construction.
Let $d,n,m$ be positive integers, $n>d,m>d$, and $j$ an integer such that
$0\leq j\leq n+m-d.$
Given two Bruhat orders $\prec$ for (${[n]}\choose{d}$ and $\prec^{\prime}$ for (${[m]}\choose{d}$ we will define a new Bruhat order $\prec\circ_{j}\prec^{\prime}$ for (${[n+m-d]}\choose{d}$, to be called the $j$-th insertion of a $\prec^{\prime}$ into $\prec$.
5.6.1 Example
If $m=2d$ then $m+n-d=n+d$. $\Box$
We define such an insertion by specifying its set of inversions.
Namely we define a subset $I(\prec\circ_{j}\prec^{\prime})\subset$ (${[n+m-d]}\choose{d+1}$ and afterwards prove that it satisfies (Z), whence it uniquely defines an order $\prec\circ_{j}\prec^{\prime}$.
Let $Inv(\prec)$ and $Inv(\prec^{\prime})$ be the sets of inversions of $\prec$ and $\prec^{\prime}$, respectively.
Consider a decomposition
$[n+m-d]=[j]\cup(j+[m])\cup(j+m+[n-(j+d)]).$
We denote $A:=[j]$, $B:=j+[m]$, $D=j+[d]\subset B$, and $C:=j+m+[n-(j+d)]$.
(i) First, for a $d+1$ element subset $J=\{j_{1},\cdots,j_{d+1}\}\subset B$, we say that
$J\in I(\prec\circ_{j}\prec^{\prime})\mbox{ iff }J-j=\{j_{1}-j,\cdots,j_{d+1}-j\}\in Inv(\prec^{\prime}).$ (5.1)
(ii) Second, for a $d+1$ element subset
$I=\{i_{1},\cdots,i_{d+1}\}\subset A\cup D\cup C$
we say that
$I\in I(\prec\circ_{j}\prec^{\prime})\mbox{ iff }\phi^{-1}(I)\in Inv(\prec),$ (5.2)
where
$\phi:[n]\to A\cup D\cup C$
is the unique monotone bijection.
(iii) Third, let $L=\{l_{1},\cdots,l_{d+1}\}\subset[n+m-d]$ be none of types (i) or (ii). Denote
$L^{+}=L\cap(A\cup D\cup C)$
and
$L^{-}=L\cap(B\setminus D)=L\setminus L^{+}.$
Thus $L=L^{-}\coprod L^{+}$.
Let
$f:L^{-}\to D$
be the unique monotone map sending $L^{-}$ to the rightmost elements of $D\setminus L^{+}$.
Denote by
$\overline{L}=L^{+}\cup f(L^{-})\subset A\cup D\cup C$
(the union is disjoint). Then we say that
$L\in I(\prec\circ_{j}\prec^{\prime})\mbox{ iff }\phi^{-1}(\overline{L})\in Inv(\prec).$ (5.3)
This three step construction defines a set $I(\prec\circ_{j}\prec^{\prime})$.
Example. If $L\cap(j+[m])=\{a,b\}$, $a\leq j+d$, $b>j+d$, then
$\overline{L}=\begin{cases}L\setminus b\cup\{j+d\}&if\quad a\neq j+d\cr L\setminus b\cup\{j+d-1\}&if\quad a=j+d.\end{cases}$
$\Box$
By definition
$I(\prec\circ_{j}\prec^{\prime})=I_{1}\coprod I_{2}\coprod I_{3}$ (5.4)
where $I_{1}$ (resp. $I_{2}$, $I_{3}$) are the subsets of elements satisfying (5.1) (resp. (5.2), (5.3)).
Proposition 5.2
The set $I(\prec\circ_{j}\prec^{\prime})$ defined above satisfies (Z), whence $I(\prec\circ_{j}\prec^{\prime})=Inv(\prec\circ_{j}\prec^{\prime})$ for the uniquely defined $\prec\circ_{j}\prec^{\prime}\in B(n+m-d,d)$.
This will be proven below, see 5.12.
5.7 Example for $d=2$.
The picture below gives an example of an insertion.
We start with two elements $w_{1}\in B(5,2)$ drawn as a collection of $5$ wires, and $w_{2}\in B(4,2)$ drawn as a collection of $4$ wires.
In the picture below the element $w_{1}\circ_{2}w_{2}\in B(7,2)$ is drawn.
Corollary 5.3
Let $\prec^{\prime}$ be dominated by $\prec^{\prime\prime}$, i.e.
$Inv(\prec^{\prime})\subset Inv(\prec^{\prime\prime})$. Then $Inv(\prec\circ_{j}\prec^{\prime})\subset Inv(\prec\circ_{j}\prec^{\prime\prime})$.
Proof. Let
$Inv(\prec\circ_{j}\prec^{\prime\prime})=J_{1}\coprod J_{2}\coprod J_{3}$
be the decomposition (5.4) for $Inv(\prec\circ_{j}\prec^{\prime\prime})$. Then $I_{1}\subset J_{1}$, whereas $I_{2}=J_{2}$ and $I_{3}=J_{3}$. $\Box$
Corollary 5.4
Let $\prec$ be dominated by $\widetilde{\prec}$, i.e.
$Inv(\prec)\subset Inv(\widetilde{\prec})$. Then $Inv(\prec\circ_{j}\prec^{\prime})\subset Inv(\widetilde{\prec}\circ_{j}\prec^{\prime})$.
Proof similar to the proof of the previous corollary.
5.8 Proof of Thm 4.3
follows from the above two corollaries since the map $\gamma$ is a composition of insertions. $\Box$
5.9 Weak Bruhat orders on symmetric groups
For a symmetric group $S(k)$, a subset $W\subset\Omega:=\{(i,j),i<j,i,j\in[k]\}$ is an inversion set of a permutation of $S(k)$ if and only if for any $i<j<l$ if $(i,j)$ and $(j,l)\in W$, then $(i,l)\in W$, and if $(i,l)\in W$, then at least one of $(i,j)$ and $(j,l)\in W$. This is nothing but Theorem 5.1.
Recall that the weak Bruhat order $\tau\prec\tau^{\prime}$ is defined by the inclusion $Inv(\tau)\subset Inv(\tau^{\prime})$.
A packet is a pair $(i,j)$, $i<j$ and a weak Bruhat order is a linear order on $[n]$ such that for each pair $(i,j)$ either $i\prec j$, or $i\succ j$. This is nothing but an aritrary permutation of $[n]$.
5.10 Insertion of permutations.
Let $\sigma\in S(n),\tau\in S(m)$, $1\leq j\leq n$. Let us define $\sigma\circ_{j}\tau\in S_{n+m-1}$ as follows.
We will use two decompositions
$[n]=[j-1]\cup\{j\}\cup(j+[n-j])=A\cup\{j\}\cup B$
and
$[n]=[\sigma(j)-1]\cup\{\sigma(j)\}\cup(\sigma(j)+[n-\sigma(j)])=A^{\prime}\cup\{\sigma(j)\}\cup B^{\prime}.$
Obviously $\sigma$ induces an isomorphism
$\sigma:A\cup B\overset{\sim}{\to}A^{\prime}\cup B^{\prime}.$
Denote for brevity $j^{\prime}:=\sigma(j)$.
Now let us stretch our decompositions to
$[n+m-1]=A\cup j-1+[m]\cup m+B$
and
$[n+m-1]=A^{\prime}\cup j^{\prime}-1+[m]\cup m+B^{\prime}.$
As before, $\sigma$ induces an isomorphism
$\sigma:A\cup m+B\overset{\sim}{\to}A^{\prime}\cup m+B^{\prime}.$
Moreover, $\tau$ induces an isomorphism
$\tau:j-1+[m]\overset{\sim}{\to}j^{\prime}-1+[m].$
Now we define $\rho:=\sigma\circ_{j}\tau\in S(n+m-1)$ as follows:
$\rho(x)=\sigma(x)$ if $x\in A\cup m+B$,
$\rho(x)=\tau(x)$ if $x\in j-1+[m]$.
5.11 Proof of Proposition 5.2 for $d=1$.
For weak Bruhat orders $\prec$ and $\prec^{\prime}$ on $[n]$ and $[m]$, let $\tau$ and $\tau^{\prime}$ be the corresponding permutations.
Due to the above construction, for $j\in[n]$, the set $Inv(\prec\circ_{j}\prec^{\prime})$ consists of pairs $(j+l,j+k)$, $l<k\leq j+m$, if $(l,k)\in Inv(\tau^{\prime})$ or pairs $(l,j+s)$, $l<j$, $s\leq m$, if $(l,j)\in Inv(\tau)$, or $(j+t,k)$, $t\leq m$, $j+t<k$ if $(j,k-m+1)\in Inv(\tau)$.
But such a set of inversions is exactly the inversion set of $\tau\circ_{j}\tau^{\prime}$. $\Box$
5.12 Proof of Proposition 5.2: arbitrary $d$.
We demonstrate that the Ziegler criterion holds for the above defined set $I(\prec\circ_{j}\prec^{\prime})$.
Let $L=\{l_{1},\cdots,l_{d+2}\}\subset[n+m-d]$, $l_{1}<\ldots<l_{d+2}$. We use the notations of 5.6.
(i) If $L\subset B$ then the Ziegler condition for the packet $P(L)$ holds true due to (5.1).
(ii) If $L\subset A\cup D\cup C$ then the Ziegler condition for the packet $P(L)$ holds true because of (5.2).
(iii) Now let $L$ be not of the above two types. We define as before
$L^{+}=L\cap(A\cup D\cup C),$
$L^{-}=L\setminus L^{+}=L\cap(B\setminus D).$
Thus $L=L^{+}\cup L^{-}$.
The packet $P(L)$ takes the form
$P(L)=\{L\setminus l_{d+2},\cdots,L\setminus l_{1}\}=\{L_{d+2},\ldots,L_{1}\}$
(iii-a) Consider first the case $L\cap B=\{l_{2},\cdots,l_{d+2}\}$.
Due to (5.3)
In such a case
$\{\phi^{-1}(L_{d+2}),\ldots,\phi^{-1}(L_{1})\}=$
$\{(l_{1},j+1,\cdots,j+d),\ldots,(l_{1},j+1,\cdots,j+d),L_{1}\}.$
Obviously, $P(L)$ satisfies the Ziegler condition whether or not
$(l_{1},j+1,\cdots,j+d)$ belongs to $Inv(\prec)$ and $L_{1}$ belongs to $Inv(\prec^{\prime})$.
(iii-b)
Let
$g:\ L^{-}\to D$
be the unique monotone map sending $L^{-}$ to the rightmost elements of $D\setminus L^{+}$. Denote by
$\overline{L}=L^{+}\cup g(L^{-})\subset A\cup D\cup C$
(the union is disjoint).
For $L\cap C=\{l_{t+1},\cdots,l_{d+2}\}$, there holds
$\overline{L\setminus l_{j}}=\overline{L}\setminus l_{j},\quad j=t+1,\cdots,d+2.$
For $L^{-}=\{l_{s+1},\cdots,l_{t}\}$, there holds
$\overline{L\setminus l_{j}}=\overline{L}\setminus g(l_{s+1}),\quad j=s+1,\cdots,t.$ (5.5)
For $L\cap(A\cup D)=\{l_{r+1},\cdots,l_{s}\}$, we have
$\overline{L\setminus l_{j}}=\begin{cases}\overline{L}\setminus g(l_{s+1}),&if\quad l_{j}>g(l_{s+1})\cr\overline{L}\setminus l_{j},&if\quad l_{j}<g(l_{s+1}).\end{cases}$
For $L\cap A=\{l_{1},\cdots,l_{r}\}$, there holds
$\overline{L\setminus l_{j}}=\overline{L}\setminus l_{j},\quad j=1,\cdots,r.$
Thus, for such $L$, the sets of the packet $P(L)$ coincide with the sets of the packet $P(\overline{L})$ except for $(t-s)$ sets corresponding to deleting from $\overline{L}$ elements of $g(L^{-})$. Namely, sets of $P(L)$ corresponding to deleting elements of $L^{-}$ from $L$, are the same, according to (5.5), and equal to
$\overline{L}\setminus g(l_{s+1}).$
However, $\overline{L}\setminus g(l_{s+1})$ is a set of the packet $P(\overline{L})$ and the packet $P(L)$, and sets of $P(\overline{L})$ are the same at the interval ending interval of $P(L)$ which starts at $\overline{L}\setminus g(l_{s+1})$, and at the beginning interval which ends on the set obtained by deleting $l_{t+1}$. Therefore, the Ziegler condition for $P(L)$ holds true, since the beginning or ending interval of inversion sets of $P(\overline{L})$ gives also the beginning or the ending interval of inversion sets of $P(\overline{L})$, and the difference is the repeating set $\overline{L}\setminus g(l_{s+1})$ in $P(L)$.
$\Box$
More briefly:
(iii-a) If $d+1$ elements of $L$ are in $B$ then the packet $P(L)$ has a form
$P(L)=(M,\ldots,M,L_{1}).$
Such a packet obviously satisfies (Z).
(iii-b) There are at most $d$ elements of $L$ in $B$.
Let
$g:\ L^{-}\to D$
be the unique monotone map sending $L^{-}$ to the rightmost elements of $D\setminus L^{+}$. Denote by
$\overline{L}=L^{+}\cup g(L^{-})\subset A\cup D\cup C$
(the union is disjoint).
The packet $P(L)$ has a form
$P(L)=(\overline{L}_{d+2},\cdots,\overline{L}_{t+1},\overline{L}_{t},\overline{L}_{t},\ldots,\overline{L}_{t},\overline{L}_{t-s},\overline{L}_{t-s-1},\ldots,\overline{L}_{1})$
where $s=Card(L^{-}\setminus D)$. Obviously such a collection satisfies the Ziegler criterion if the packet $P(\overline{L})$ satisfies it. $\Box$
5.13 Associativity and commutativity
Lemma 5.5
The insertion $\prec\circ_{j}\prec^{\prime}$ is associative. That is
$(\prec\circ_{j}\prec^{\prime})\circ_{j+j^{\prime}}\prec^{\prime\prime}=\prec\circ_{j}(\prec^{\prime}\circ_{j^{\prime}}\prec^{\prime\prime}).$
Proof. Consider a $d+1$-element subset $L$ of $[n+m+m^{\prime}-2d]$ obtained by insertion of $[m]$ instead of the team $j+[d]$ into $[n]$ and $[m^{\prime}]$ instead of the team $j+j^{\prime}+[d]$ into $[n+m-d]$. If $L$ has a non-empty intersection with $A$ or $C$, then (5.2) and (5.3) gives the same $\overline{L}$ for $(\prec\circ_{j}\prec^{\prime})\circ_{j+j^{\prime}}\prec^{\prime\prime}$ and $\prec\circ_{j}(\prec^{\prime}\circ_{j^{\prime}}\prec^{\prime\prime})$.
If $L\subset(j+[m])$ then by (5.1), $\overline{L}$ is defined by (5.1-5.3) for $\prec^{\prime}\circ_{j^{\prime}}\prec^{\prime\prime}$, that is transformed following (5.1) or (5.3). Both transformation gives the same change of $L$ for $(\prec\circ_{j}\prec^{\prime})\circ_{j+j^{\prime}}\prec^{\prime\prime}$, since $L$ does not change in $\prec\circ_{j}\prec^{\prime}$, that is it is either inverse for both $\prec\circ_{j}\prec^{\prime}$ and shifted $\prec^{\prime}$ or not. $\Box$
The same arguments give the commutativity property.
Lemma 5.6
For a positive integer $d$ and nonnegative integers $k_{0},k_{1}$,
there holds
$(\prec\circ_{k_{0}}\prec^{\prime})\circ_{k_{0}+m+k_{1}}\prec^{\prime\prime}=(\prec\circ_{k_{0}+d+k_{1}}\prec^{\prime\prime})\circ_{k_{0}}\prec^{\prime}.$
$\Box$
CHECK THE INDICES!
6 Operads of higher Bruhat orders
6.1 Two operads
Fix a positive integer $d$. In this Section we will define two planar operads. The first one, to be denoted $HB_{0}^{d}$ and called the small ($d$-th) Bruhat operad, and the second one, to be denoted $HB^{d}$ and called the big ($d$-th) Bruhat operad.
The operad $HB_{0}^{d}$ will be a suboperad of $HB^{d}$.
6.2 Small Bruhat operad.
Denote $HB_{0}^{d}(n):=B(nd,d)$.
Define composition maps
$\gamma:HB_{0}^{d}(n)\times HB_{0}^{d}(m_{1})\times\cdots\times HB_{0}^{d}(m_{n})\to HB_{0}^{d}(m_{1}+\cdots+m_{n})$ (6.1)
by the rule
$\gamma(b_{0};b_{1},\ldots,b_{n})=$
$=(((b_{0}\circ_{0}b_{1})\circ_{dm_{1}}\cdots)\circ_{dm_{1}\cdots+dm_{n-1}}b_{n}).$
Theorem 6.1
The collection $HB_{0}^{d}(n)$, $n\geq 1$, with the compositions (6.1) forms a planar operad $HB_{0}^{d}$.
Proof. Follows from Proposition 5.2, Lemmata 5.5, 5.6, and 2.3. $\Box$
6.3 Big Bruhat operad.
Denote by $HB^{d}(n)$ an infinite set whose elements are triples
$\alpha=(m,b,\mathbf{k})$
where $m$ is an integer $m>d$, $b\in B(m,d)$, and $\mathbf{k}$ is a type with $|\mathbf{k}|=m$ such that the molecule $M(\mathbf{k})$ has $n$ nuclei, each of them containing $d$ protons.
For notations see 3.2.
Define composition maps
$\gamma:HB^{d}(n)\times HB^{d}(m_{1})\times\cdots\times HB^{d}(m_{n})\to HB^{d}(m_{1}+\cdots+m_{n})$ (6.2)
by the following rule. Let $\alpha_{0}=(m_{0},b_{0},\mathbf{k}_{0})\in HB^{d}(n)$, $\alpha_{i}=(m^{\prime}_{i},b_{i},\mathbf{k}_{i})\in HB^{d}(m_{i}),1\leq i\leq n$. Let
$\mathbf{k}_{0}=(k_{0},d,k_{1},\ldots,d,k_{n})$
Then
$\gamma(\alpha_{0};\alpha_{1},\ldots,\alpha_{n})=$
$=(((b_{0}\circ_{k_{0}}b_{1})\cdots)\circ_{k_{0}+m_{1}+k_{1}+\cdots+k_{n-2}+m_{n-1}}b_{n}).$
Theorem 6.2
The collection $HB^{d}(n)$ $n\geq 1$ with the compositions (6.2) forms a planar operad $HB^{d}$. We have an embedding of operads $HB_{0}^{d}\subset HB^{d}$.
Proof. Follows from Proposition 5.2, Lemmata 5.5, 5.6, and 2.3. $\Box$
Due to Corollary 5.3 these planar operads are compatible with the $d$-th Bruhat order in the same sense as the symmetric group operad $F$ is compatible with the weak Bruhat order on symmetric groups.